OpenStudy (anonymous):

stephanie has the following scores on her algebra tests: 85, 92, and 71. What score must she earn on her fourth test in order to have an average of 80 or above?

OpenStudy (anonymous):

Finding the average of a set of data is found by adding up all the data and dividing it by the amount of data that you have. In this case, that would be: (85 + 92 + 71 + x) ÷ 4 = 80 x = The minimum score Stephanie needs to save her from getting a C in the algebra class. From there, it's just a solve for x problem. (85 + 92 + 71 + x) ÷ 4 = 80 Multuply each side by 4. (85 + 92 + 71 + x) = 320 Simplify the left side. 248 + x = 320 Subtract 248 from both sides. x = 72 So, Stephanie needs at least a 72 on her fourth test to get a B in the class. Knowing her previous scores, it shouldn't be that hard as long as it's nothing like her third test.

OpenStudy (anonymous):

lol tks alot

OpenStudy (anonymous):

No problem. :D